OpenStudy (anonymous):

what is the reference angle to 47pi/6 in radians

OpenStudy (tkhunny):

Hints: 47 - 12 = 35 35 - 12 = 23 23 - 12 = 11 11 - 12 = -1

OpenStudy (anonymous):

pi over 6